At a certain temperature and pressure, one liter of CO2 gas weighs 1.85 g. What is the mass of one liter of NH3 gas at the same temperature and pressure?
Answer #1

Since P,T is same for both. Even volume is same which is 1 L

number of mol should also be same for both

Molar mass of CO2,

MM = 1*MM(C) + 2*MM(O)

= 1*12.01 + 2*16.0

= 44.01 g/mol

mass(CO2)= 1.85 g

number of mol of CO2,

n = mass of CO2/molar mass of CO2

=(1.85 g)/(44.01 g/mol)

= 4.204*10^-2 mol

number of moles of NH3 should also be 4.204*10^-2 mol

Molar mass of NH3,

MM = 1*MM(N) + 3*MM(H)

= 1*14.01 + 3*1.008

= 17.034 g/mol

mass of NH3,

m = number of mol * molar mass

= 4.204*10^-2 mol * 17.034 g/mol

= 0.716 g

Answer: 0.716 g